OpenStudy (anonymous):

Solve for x: (5)/(x^2-4) + (2)/(x) = (2)/(x-2)

OpenStudy (anonymous):

\[\frac{ 5 }{ x ^{2}-4 } + \frac{ 2 }{ x-2}\]

OpenStudy (tyteen4a03):

Remember that \(x^2 - 4 = x^2 - 2^2\).